OpenStudy (anonymous):

In a large lecture class, there were 53 more men than women. If there were 329 students in the class, how many were in the class......

OpenStudy (lgbasallote):

let m = number of male and w = women it says there were 53 more men than women so m = 53 + w does that help?

OpenStudy (smurfy14):

329=w+w+53 276=2w woment=138

OpenStudy (anonymous):

what about 53 (men) + x = 329

OpenStudy (smurfy14):

and men would be whatever 329-138 is

OpenStudy (anonymous):

still confused

OpenStudy (lgbasallote):

wait...if your question is "how many were in the class" wouldn't it just be 329 because you said there were 329 students in class?

OpenStudy (anonymous):

if there were 329 students in the class, how many women were in the class?